English: Solve this integral
PARTIE B: QUESTIONS À DÉVELOPPEMENT : Question 10. [5 points) Calculer lintégrale indéfinie suivante.
Answer #1

We evaluate the value of the given indefinite integral by using partial fraction decomposition method. Given infegral l 23 + n du 2-2011 : Now = 2(x+) du partial fraction decomposition method . using seta (x) ant at a lat) + MaAgain put neo in U and get 0 (04+0) =0(0-1)+ A (0-1)+ B(0-1)+c a oz 0+ A -B+ 2 [::(=2] .* . B-A22 .. (2) again put na 2 inubito Therefore the given infegral can be written a sering dan floring treating de In + A+ m 3nt Ju du + 3 aut sfat, da tz in